OpenStudy (anonymous):

what are the coordinates ofthe point halfway between (-1,1) and (3,1)?

OpenStudy (anonymous):

(1, 1)

OpenStudy (anonymous):

coordinates ofthe point halfway = \[(\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2}) = (\frac{-1 + 3}{2} , \frac{1 + 1}{2}) \] \[= (\frac{2}{2} , \frac{2}{2}) = (1, 1)\] @fanman